Circular No. 7834

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS K. Hinkle, R. Joyce, and L. Wallace, National Optical Astronomy Ob-

servatory; and B. Rodgers, Gemini Observatory, report high-resolution (R = 50 000) infrared spectroscopy of the 2.3-μm region of V838 Mon, obtained with the Phoenix spectrograph on the Gemini South telescope on Feb. 13.1 UT: “CO appears to have weakened dramatically from previous reports (IAUC 7796 and 7829), with absorption central depths of at most 4 percent for individual 2-0 R-branch lines. The 3-1 band head is not prominent. The CO lines are broad with FWMH = 50 km/s at a heliocentic radial velocity of −10 km/s. The CO excitation temperature is that of a stellar photosphere. A few atomic lines are present in absorption in the region observed. Na I lines at 2336 and 2338 nm have a velocity of +190 km/s, confirming the presence of multiple-velocity systems in V838 Mon (IAUC 7786, 7812, 7822, 7829). Using Phoenix in imaging mode, the [2320 nm] magnitude was determined to be +4.1. This implies significant brightening in the infrared from mid-January (IAUC 7791 and 7796).”

SUPERNOVA 2002ap IN M74 K. Motohara and K. Nomoto, University of Tokyo; C. Gerardy and

R. Fesen, Dartmouth College; and P. Henry, University of Hawaii, report on near-infrared imaging and spectroscopy of SN 2002ap with the 8.3-m Subaru telescope (+ CISCO) on Feb. 7.21–7.26 UT. The spectrum (range 0.88–2.5 μm, resolution 220–440) shows a strong P-Cyg emission feature with a broad absorption minimum near 1.03 μm, along with weaker features with minima near 0.96, 1.51, and 2.01 μm. If the strong 1.03-μm feature is attributed to He I (1.083 μm), then the observed blueshift implies a photospheric expansion velocity of roughly 16 000 km/s. Mg II 1.093-μm and C I 1.069-μm could also be contributing to this feature. Estimated infrared magnitudes: Feb. 7.214, J = 11.97 ± 0.05; 7.217, H = 11.77 ± 0.05; 7.219, K = 11.68 ± 0.05.

SUPERNOVA 2002aw IN ANONYMOUS GALAXY T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002aw (cf. IAUC 7831), obtained by M. Calkins on Feb. 18.52 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. The spectral-feature age (cf. IAUC 7774) of the supernova is 1 ± 2 days before maximum light.

2002 February 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green